OpenStudy (anonymous):

factor fully: 2xy+4x+3y+6 9x^2-24x+16

OpenStudy (anonymous):

sorry those are supposed to be seperate: \[2xy+4x+3y+6\] \[9x^{2}-24x+16\]

OpenStudy (anonymous):

2x(y+2)+3(y+2) (y+2)(2x+3)

OpenStudy (anonymous):

9x^2-12x-12x+16 Use the same principle

OpenStudy (anonymous):

2xy+4x+3y+6 = 2x(y +2) + 3(y +2) = (2x +3) (y + 2)

OpenStudy (anonymous):

9x2−24x+16 = ( 3x - 4 )^2 = 0 => x = 4/3